Is the a true or false information and please show your work

10.00g calcium carbonate is needed to prepare 1.0L solution such that 50.0ml of this solution diluted to 200ml gives a .5%w/v solution
Answer #1

Sol . Information given is false .

Explanation :

Firstly , we have to find the concentration of concentrated solution .

As C1 V 1 = C2 V2

where C1 and C2 are the concentrations for concentrated and diluted solutions respectively , V1 and V2 are the volumes for concentrated and diluted solutions respectively .

Now , C2 = 0.5 % w/v , V2 = 200 ml , V1 = 50 ml

So , C1 = C2 V2 / V1 = 0.5 × 200 / 50 = 2 % w/v

As 2% w / v means 2 g of calcium carbonate is required for 100 ml of solution .

So , For 1L(1000 ml ) of solution , mass of calcium carbonate required = 2×1000/100 = 20 g

Therefore ,  20.00g of calcium carbonate is needed to prepare 1.0 L of solution such that 50.0 ml of this solution diluted to 200 ml gives a 0.5 % w/v solution .